# SNOOPI: Supercharged One-step Diffusion Distillation with Proper Guidance

### v2

@@ 7 Discussion and Conclusion
Limitations: Despite its promising results, PG-SB currently lacks support for few-step models, limiting its applicability to downstream tasks. Additionally, while NASA proves effective, achieving high-quality images depends on selecting an appropriate scale for removing negative features, though the parameter does not require extensive tuning. Furthermore, NASA is designed specifically for architectures with cross-attention layers, leaving support for other architectures, such as MM-DiT [9], as future work.
Conclusion: This paper introduces SNOOPI, a framework that improves stability and control in one-step diffusion models for both training and inference. PG-SB dynamically adjusts the guidance scale to stabilize VSD across various model backbones without increasing computational cost. Additionally, NASA is the first approach to incorporate negative prompt guidance in one-step diffusion models, enabling the effective reduction of unwanted features in generated images through cross-attention adjustments. Experimental results demonstrate that SNOOPI outperforms strong baselines, showcasing its stability and effectiveness.
@@ 8 Implementation Details
PG-SB: Following SBv2 [7], we also use the clamped CLIP loss with a margin of $\tau=0.37$, starting with a weight of 0.1 and gradually reducing to zero. Tab. 6 provides the additional hyperparameters for training PG-SB.
Hyperparameter SDv1.5 SDv2.1 PixArt-$\alpha$
Dataset JDB + LAION JDB + LAION JDB
Batch size 64 64 64
Training iterations 60k 40k 60k
Mixed-Precision (BF16) Yes Yes Yes
$(\kappa_{\text{min}},\kappa_{\text{max})}$ (0.5, 4) (0.5, 4) (0.5, 3)
Clip weight 0.1 0.1 0.1
$\tau$ 0.37 0.37 0.37
$lr$ of student $1e^{-6}$ $1e^{-6}$ $1e^{-6}$
$lr$ of LoRA teacher $1e^{-3}$ $1e^{-3}$ $1e^{-3}$
LoRA rank $r$ 64 64 64
LoRA scaling $\gamma$ 128 128 128
Table 6: Hyperparameters used for training PG-SB.
Positive prompt Negative prompt SDXL-LCM SDXL-DMD2 SDXL-DMD2 PG-SB
4 steps 4 steps 1 step 1 step
“A photo of a person” “male” 52% $\rightarrow$ 99% 26% $\rightarrow$ 100% 29% $\rightarrow$ 100% 42% $\rightarrow$ 81%
“female” 48% $\rightarrow$ 100% 74% $\rightarrow$ 100% 71% $\rightarrow$ 100% 58% $\rightarrow$ 88%
“A photo of a person” “young” 55% $\rightarrow$ 100% 44% $\rightarrow$ 100% 7% $\rightarrow$ 100% 39% $\rightarrow$ 96%
“old” 45% $\rightarrow$ 92% 56% $\rightarrow$ 100% 93% $\rightarrow$ 100% 61% $\rightarrow$ 79%
“A photo of a pet” “cat” 36% $\rightarrow$ 100% 12% $\rightarrow$ 100% 60% $\rightarrow$ 100% 32% $\rightarrow$ 99%
“dog” 64% $\rightarrow$ 100% 88% $\rightarrow$ 100% 40% $\rightarrow$ 100% 68% $\rightarrow$ 99%
Table 7: Comparison of success rates of unwanted feature removal in generated images before and after applying the NASA model.
Method Anime Photo Concept Art Paintings Average
Stable Diffusion 1.5-based backbone
InstaFlow-0.9B [26] 26.10 26.62 25.92 25.95 26.15
InstaFlow-0.9B + NASA ($\alpha=0.19$) 26.24 26.74 26.04 26.00 26.26 (+0.11)
DMD2 [54] 26.39 27.00 25.80 25.83 26.26
DMD2 + NASA ($\alpha=0.04$) 26.41 27.02 25.80 25.83 26.27 (+0.01)
PG-SB 27.18 27.58 26.69 26.62 27.02
PG-SB + NASA ($\alpha=0.04$) 27.19 27.59 26.71 26.63 27.03 (+0.01)
Stable Diffusion 2.1-based backbone
SBv2 [7] 27.25 27.62 26.86 26.77 27.13
SBv2 + NASA ($\alpha=0.30$) 27.45 27.85 26.93 27.09 27.33 (+0.20)
PG-SB 27.56 27.84 26.97 27.03 27.35
PG-SB + NASA ($\alpha=0.30$) 27.71 27.99 27.14 27.27 27.53 (+0.18)
PixArt-$\alpha$-based backbone
YOSO [30] 28.79 28.09 28.57 28.55 28.50
YOSO + NASA ($\alpha=0.20$) 28.80 28.10 28.62 28.57 28.52 (+0.02)
DMD [55] 29.31 28.67 28.46 28.41 28.71
DMD + NASA ($\alpha=0.45$) 29.34 28.71 28.50 28.52 28.77 (+0.06)
PG-SB 32.19 29.09 30.39 29.69 30.34
PG-SB + NASA ($\alpha=0.80$) 32.56 29.55 31.24 30.96 31.08 (+0.74)
Table 8: HPSv2 comparisons between our method and previous works. denotes reported numbers, denotes our rerun based on the publicly available model checkpoints.
NASA: We provide the PyTorch pseudo-code for the NASA Attention algorithm, as outlined in Sec. 5. The implementation is straightforward and is provided in Algorithm 1.
Algorithm 1 NASA Attention.
class NASA_AttnProcessor(nn.Module):
def __init__(self, scale=1.0):
@@
self.scale = scale
@@
self, attn, z, emb, neg_emb, attn_mask, neg_attn_mask, temb, scale=1., *args, **kwargs,
@@
z_nasa = z - self.scale * neg_z
@@
9 More Analysis on NASA
9.1 Experimental Details in Tab. 3
For clarification, in Tab. 3, we evaluate the effectiveness of NASA across four models: SDXL-LCM, SDXL-DMD2 (with both 1-step and 4-step settings), and our PG-SB model using the SDv2.1 backbone. For the “without NASA” case, all models use a scale of $\alpha=0$, while for the “with NASA” case, a scale of $\alpha=0.5$ is applied. The evaluation involves 6 positive-negative prompt pairs. For each model and prompt pair, 100 images are generated. Human evaluations are then used to determine the percentage of images that successfully exclude the feature specified by the negative prompt. The detailed results from Tab. 3 are provided in Tab. 7.
9.2 More Quantitative Results
Tab. 8 presents additional quantitative results when applying NASA to various models. Using a negative prompt such as “worst quality, low quality, ugly, duplicate, out of frame, deformed, blurry, bad anatomy, watermark,” our NASA method demonstrates a slight improvement in the HPSv2 scores across all models.
10 Additional Qualitative Results of PG-SB
Fig. 7, Fig. 8, and Fig. 9 present additional uncurated samples synthesized by our model using three different backbones: PixArt-$\alpha$, SDv1.5, and SDv2.1, respectively.
Figure 7: Additional qualitative images generated by our PG-SB model with the PixArt-$\alpha$ backbone.
Figure 8: Additional qualitative images generated by our PG-SB model with the SDv1.5 backbone.
Figure 9: Additional qualitative images generated by our PG-SB model with the SDv2.1 backbone.

### v3

@@ 7 Discussion and Conclusion
This paper introduces NASA, the first method to integrate negative prompt guidance into one-step and few-step diffusion models for both inference and distillation. By leveraging cross-attention adjustments, NASA effectively suppresses unwanted features in generated images. Experimental results show that NASA enhances image quality and outperforms strong baselines, establishing a new state-of-the-art in one-step model distillation based on human preference metrics. While NASA’s effectiveness relies on selecting an appropriate scale for suppressing negative features, this parameter does not require extensive tuning.
@@ 8 Implementation Details
NASA. We provide the PyTorch pseudo-code for the NASA algorithm, as outlined in Section 5.1 in the main paper. The implementation is straightforward and is provided in Algorithm 1.
Algorithm 1 NASA.
class NASA_AttnProcessor($n$.Module):
def __init__(self, nasa_scale=1.0):
@@
self.nasa_scale = nasa_scale
@@
self, attn, z, emb, neg_emb, attn_mask, neg_attn_mask, temb, *args, **kwargs,
@@
z_nasa = z - self.nasa_scale * neg_z
@@
NASA–T. Following SBv2 [13], we also use the clamped CLIP loss with a margin of $\tau=0.37$, starting with a weight of 0.1 and gradually reducing to zero. Table 6 provides the additional hyperparameters for training SBv2 with NASAT.
Hyperparameter SDv1.5 SDv2.1 PixArt-$\alpha$
Dataset JDB + LAION JDB + LAION JDB
Batch size 64 64 32
Training iterations 60k 40k 50k
Mixed-Precision (BF16) Yes Yes Yes
$\kappa$ $\mathcal{U}(0.5,4)$ $\mathcal{U}(0.5,4)$ $\mathcal{U}(0.5,3)$
$\alpha$ $\mathcal{U}(0,1)$ $\mathcal{U}(0,1)$ $\mathcal{U}(0,1)$
Clip weight 0.1 0.1 0.1
$\tau$ 0.37 0.37 0.37
lrlr of student $1e^{-6}$ $1e^{-6}$ $1e^{-6}$
lrlr of LoRA teacher $1e^{-3}$ $1e^{-3}$ $1e^{-3}$
LoRA rank $r$ 64 64 64
LoRA scaling $\gamma$ 128 128 128
Table 6: Hyperparameters used for training SBv2 [13] with NASA–T.
9 More Qualitative Results
Figure 8 and Figure 9 illustrate the visual effect of different scale values $\alpha$ ranging from 0.0 to 1.0 in the NASA method when being applied on SDXL-DMD2. A higher scale value $\alpha$ results in more effective removal of the feature specified by the negative prompt.
Figure 8: Additional qualitative images of applying NASA–I to SDXL-DMD2 1-step.
Figure 9: Additional qualitative images of applying NASA–I to SDXL-DMD2 4-step.